A 5kg object which is initially at rest is acted upon by a 25N net external force. What is the speed of the object after 4 seconds?
1
Expert's answer
2019-10-22T10:19:38-0400
"a=\frac{F}{m}=\frac{25}{5}=5\frac{m}{s^2}"

"v=at=(5)(4)=20\frac{m}{s}"